Given the following equation, C2H5I (g) ⇒ C2H4 (g) + HI (g) k= 4.0 * 10^-3 s^ -1 at 700K

a. Write the rate law.

b. what is the rate of reaction when {C2H5I}=3.3x10^-4mol/L?

c. if the concentration is doubled, what is the effect on the reaction?

d. What is the half life?

2. Given the following equation 2C2H2=C4H4 K=0.178 L/mol at 770K

a. write the rate law

b. what is the rate of reaction when C2H2=0.195M

c. what is the half life

d. How much is required for 70% of the C2H2 to be used up?
Answer #1

1.

(a)

Rate law,

Rate = k[C2H5I]

From the units of k , it is understood that the reaction is first order reaction.

(b)

Rate = 4.0 * 10-3 * 3.3 * 10-4

Rate = 1.32 * 10-6 M/s

(c)

If the concentration is doubled the rate of reaction is also doubled.

(d)

Half life is the time required for the conversion of half of the reactant into product.

Half life time = 0.693 / k = 0.693 / (4.0 * 10-3) = 173. s

2.

(a)

From the units of k , it is understood that the reaction is second order reaction.

Rate = k[C2H2]2

(b)

Rate = 0.178 * (0.195)2 = 6.77 * 10-3 M/s

(c)

Half life time = 1 / k[A]0 = 1 / (0.178 * 0.195) = 28.8 s

(d)

Initial [A]0 = 0.195 M

Final [A] = 0.195 - (0.195 * 70 / 100) = 0.0585 M

Second order rate constant equation is,

1 / [A] = kt + 1 / [A]0

1 / 0.0585 = (0.178 * t) + ( 1 / 0.195)

t = 67.2 s